OpenStudy (anonymous):

The point (4,-4)would not fall on the line for which of the following representations 2Y= -3X + 4 7 - 3Y=4X + 3 18 - 6X=3Y + 6 -4Y = 6X + 8

OpenStudy (ash2326):

@KALYNNKANE here we've to check each of the options and find the equation of the line which doesn't have the point (4, -4) on it I'll check first option for you (4, -4) x=4 y=-4 \[2Y=-3X+4\] \[2\times -4=-3\times 4+4\] \[-8=-12+4\]\[-8=-8\] this is true, so this is not our answer. Would you check the other options?

OpenStudy (anonymous):

THEN ITS D

OpenStudy (ash2326):

Good:) did you understand this?

OpenStudy (anonymous):

LITTLE

OpenStudy (ash2326):

Where do you have confusion?

OpenStudy (anonymous):

THE ANSWER LOL

OpenStudy (ash2326):

did you check each option?

OpenStudy (anonymous):

YUP

OpenStudy (ash2326):

Did you find that the point lie on the 4th equation?

OpenStudy (anonymous):

no i didnt get the problem

OpenStudy (ash2326):

@KALYNNKANE Please read my first post, and let me know where you have confusion

OpenStudy (anonymous):

2x -4 = -3

OpenStudy (ash2326):

We have the first equation as \[2Y=-3X+4\] we have the point (4, -4) let's substitute x=4 and y=-4 here to check if the point lies on this line \[2\times -4=-3\times 3+4\]

OpenStudy (anonymous):

dont get it